# Explainable Multi-modal Time Series Prediction with LLM-in-the-Loop

### v2

@@ 1 Introduction
In the modern big-data era, time series analysis has become indispensable for understanding real-world system behaviors and guiding downstream decision-making tasks across numerous domains, including healthcare, traffic, finance, and weather [1, 2, 3, 4]. Although deep learning models have demonstrated success in capturing complex temporal dependencies [5, 6, 7, 8], real-world time series are frequently influenced by external information beyond purely temporal factors. Such additional context, which may come from textual narratives (e.g., finance news [9] or medical reports [10]), can offer critical insights for more accurate forecasting and explainability.
Recent multi-modal approaches for time series have shown promise by integrating rich contextual signals from disparate data sources—such as textual descriptions—to improve performance on tasks ranging from forecasting and classification to imputation and retrieval [11, 12, 13, 14, 15, 16, 17]. While these approaches utilize supplementary data to enhance predictive accuracy, they often lack explicit mechanisms to systematically reason and explain about why or how contextual signals affect outcomes. This gap in interpretability poses significant barriers for high-stakes applications such as finance or healthcare, where trust and transparency are paramount.
Meanwhile, Large Language Models (LLMs) [18, 19, 20] have risen to prominence for their remarkable ability to process and reason over textual data across domains, enabling tasks like sentiment analysis, question answering, and content generation in zero- and few-shot settings [21, 22, 23]. Recent advancements in agentic designs further augment LLM capabilities, facilitating structured reasoning and iterative decision-making for context-rich scenarios [24, 25, 26]. Their encoded domain knowledge makes them natural candidates for supporting real-world multi-modal time series analyses, where textual context (e.g., news or expert notes) plays a vital role [27, 28, 29, 30, 31, 32, 33].
@@
Motivated by these observations, we introduce TimeXL, a novel framework that adopts a closed-loop workflow of prediction, critique (reflect), and refinement, and unifies a prototype-driven time series encoder with LLM-based reasoning to deliver both accurate and interpretable multi-modal forecasting (Figure 1). Our approach first employs a multi-modal prototype-based encoder to generate preliminary time series predictions alongside human-readable explanations, leveraging case-based reasoning [34, 35, 36, 37] from both the temporal and textual modalities. These explanations not only justify the encoder’s predictions but also serve as auxiliary signals to guide an LLM-powered component that further refines the forecasts and contextual rationales.
@@ 2 Related Work
Our work is primarily related to three key areas of time series research: multi-modality, explanation, and the use of LLMs. A more detailed discussion of related methods is provided in Appendix G.
@@
Multi-modal Time Series Analysis. In recent years, multi-modal time series analysis has gained increasing attention in diverse domains such as finance, healthcare, environmental sciences, and industry [11, 38, 12, 39, 16, 14]. Multiple approaches have been proposed to model interactions across different modalities, including fusion and alignment through concatenation [40, 41], attention mechanism [13, 42, 43], gating functions [44], and contrastive learning [45, 46]. These methods facilitate tasks beyond forecasting and classification, such as retrieval [17], imputation [16], and causal discovery [45]. Recent research efforts have also advanced the field through comprehensive benchmarks and studies [47, 48, 49, 50, 51], highlighting the incorporation of new modalities to boost task performance. Meanwhile, others explore transforming time series to other modalities [52, 53, 54, 55] or generating synthetic ones [44, 56] for effective analysis. Nevertheless, these studies primarily focus on improving numerical performance while the deeper and tractable rationale behind why or how the textual or other contextual signals influence time series outcomes remains underexplored.
@@
Time Series Explanation. Recent studies have explored diverse paradigms for time series interpretability. Gradient-based and perturbation-based saliency methods, for example, highlight important features at different time steps [57, 58], where some works explicitly incorporate temporal structures into models and objectives [59, 60]. Surrogate approaches also offer global or local explanations, such as applying Shapley values to time series [61], enforcing model consistency via self-supervised objectives [62], or using information-theoretic strategies for coherent explanations [63]. In contrast to saliency or surrogate-based explanations, we adopt a case-based reasoning paradigm [34, 35, 36, 37], which end-to-end generates predictions and built-in explanations from learned prototypes. Our work extends this approach to multi-modal time series by producing human-readable reasoning artifacts for both the temporal and contextual modalities.
@@
LLMs for Time Series Analysis. The rapid development of LLMs [18, 19, 20] has begun to inspire new directions in time series research [64, 65]. Many existing methods fine-tune pre-trained LLMs on time series tasks and achieves promising results [66, 67, 68], where textual data (e.g., domain instructions, metadata, summaries) are often encoded as prefix embeddings to enrich time series representations [69, 70, 71, 72, 73, 74]. These techniques also contribute to the emergence of time series foundation models [68, 75, 76, 77, 78]. An alternative line of research leverages the zero-shot or few-shot capabilities of LLMs by directly prompting pre-trained language models with text-converted time series [79] or context-laden prompts representing domain knowledge [30, 32, 33], often yielding surprisingly strong performance in real-world scenarios. Furthermore, LLMs can act as knowledge inference modules, synthesizing high-level patterns or explanations that augment standard time series pipelines [80, 31, 81, 82]. Building on this trend, recent works have explored time series reasoning with LLMs by framing tasks as natural-language problems, including time series understanding & question answering [83, 84, 85, 78, 86, 87, 88, 89], and language-guided inference [83, 90, 78, 85]. Compared with existing works, our work provides a unique synergy between an explainable model and LLMs, enhancing explainable prediction through iterative, grounded, and reflective interaction.
@@ 3.1 Problem Statement
In this paper, we consider a multi-modal time series prediction problem. Each instance is represented by the multi-modal input $(\boldsymbol{x},\boldsymbol{s})$, where $\boldsymbol{x}=(x_{1},x_{2},\cdots,x_{T})\in\mathbb{R}^{N\times T}$ denotes time series data with $N$ variables and $T$ historical time steps, and $\boldsymbol{s}$ denotes the corresponding text data describing the real-world context. The text data $\boldsymbol{s}$ can be further divided into $L$ meaningful segments. Based on the historical time series and textual context, our objective is to predict the future outcome $\boldsymbol{y}$, either as a discrete value for classification tasks, or as a continuous value for regression tasks. We focus on the classification task to reflect discrete decision-making scenarios common in real-world multi-modal applications, where interpretability is often essential for reliable decision support. We also include an evaluation of the regression setting in the Appendix F. There are four major components in the proposed TimeXL framework, a multi-modal prototype encoder $\mathcal{M}_{\mathrm{enc}}$ that provides an initial prediction and case-based explanations, a prediction LLM $\mathcal{M}_{\mathrm{pred}}$ that provides a prediction based on contextual understanding with explanations, a reflection LLM $\mathcal{M}_{\mathrm{refl}}$ that generates feedback, and a refinement LLM $\mathcal{M}_{\mathrm{refine}}$ that refines the textual context based on the feedback. Below, we introduce each component and how they synergize toward better prediction and explanation.
@@ 3.2.1 Multi-modal Sequence Modeling with Prototypes
Sequence Encoder. To capture both temporal and semantic dependencies, we adopt separate encoders for time series ($\mathcal{E}_{\theta}$) and text ($\mathcal{E}_{\phi}$). For $\boldsymbol{x}\in\mathbb{R}^{N\times T}$, the time series encoder $\mathcal{E}_{\theta}$ maps the entire sequence into one or multiple representations, which serve as candidates for prototype learning. Simultaneously, the text input $\boldsymbol{s}$ is first transformed by a frozen pre-trained language model, $\mathrm{PLM}$ (e.g., BERT [91] or Sentence-BERT [92]), to produce embeddings $\boldsymbol{e}_{s}\in\mathbb{R}^{d_{s}\times L}$. These embeddings are then processed by a separate encoder $\mathcal{E}_{\phi}$ to extract meaningful text features. It is worth noting that the choice of $\mathcal{E}_{\theta}$ and $\mathcal{E}_{\phi}$ also affects the granularity of explanations. As we will introduce shortly, the prototypes are learned based on sequence representations and are associated with the counterparts in the input space, where the correspondences are determined by the encoders. In this paper, we choose convolution-based encoders for both modalities to capture the fine-grained sub-sequence (i.e., segment) patterns:
@@ 3.3.1 Model Synergy for Augmented Prediction
Note that $i,j,c$ denotes the prototype index, segment index, and class index, respectively. As $\mathbf{expl}_{\boldsymbol{s}}$ can contain relevant contextual guidance across multiple classes, it augments the input space and removes semantic ambiguity for prediction agent $\mathcal{M}_{\mathrm{pred}}$. Therefore, the prediction is drawn as $\boldsymbol{\hat{y}}_{\mathrm{LLM}}=\mathcal{M}_{\mathrm{pred}}(\boldsymbol{s},\mathbf{expl}_{\boldsymbol{s}})$. The prompt for querying $\mathcal{M}_{\mathrm{pred}}$ is provided in Appendix D, Figure 14.
@@ 3.3.2 Iterative Context Refinement via Reflective Feedback
Given the prediction $\boldsymbol{\hat{y}}_{\mathrm{LLM}}$ generated by the prediction agent $\mathcal{M}_{\mathrm{pred}}$, the reflection agent $\mathcal{M}_{\mathrm{refl}}$ aims to understand the reasoning behind the implicit prediction logic of $\mathcal{M}_{\mathrm{pred}}$. Specifically, it generates a reflective feedback, $\mathrm{Refl}$, by analyzing the input text $\boldsymbol{s}$ and its prediction $\boldsymbol{\hat{y}}_{\mathrm{LLM}}$, against the ground truth $\boldsymbol{y}$, to provide actionable insights for refinement, i.e., $\mathrm{Refl}=\mathcal{M}_{\mathrm{refl}}(\boldsymbol{y},\boldsymbol{\hat{y}}_{\mathrm{LLM}},\boldsymbol{s})$. Guided by the feedback, the refinement agent $\mathcal{M}_{\mathrm{refine}}$ refines the previous text $\boldsymbol{s}_{i}$ into $\boldsymbol{s}_{i+1}$ by selecting and emphasizing the most relevant content, ensuring that important patterns are appropriately contextualized, which is similar to how a domain expert would perform, i.e., $\boldsymbol{s_{i+1}}=\mathcal{M}_{\mathrm{refine}}(\mathrm{Refl},\boldsymbol{s}_{i})$. The prompts for querying $\mathcal{M}_{\mathrm{refl}}$ and $\mathcal{M}_{\mathrm{refine}}$ are provided in Figures 1617 1819, and discussed in Appendix D.
@@
Inputs: Multi-modal time series $(\boldsymbol{x},\boldsymbol{s},\boldsymbol{y})$ with $\mathcal{D}_{\mathrm{train}}$, $\mathcal{D}_{\mathrm{val}}$, $\mathcal{D}_{\mathrm{test}}$, prototype-based encoder $\mathcal{M}_{\mathrm{enc}}$, prediction agent $\mathcal{M}_{\mathrm{pred}}$, reflection agent $\mathcal{M}_{\mathrm{refl}}$, refinement agent $\mathcal{M}_{\mathrm{refine}}$, fusion parameter $\alpha$, max iteration $\tau$, improvement check $\mathrm{Eval(\cdot)}$ Training & Validation:
@@ 4 Experiments
DLinear [93] 0.540 0.660 0.255 0.485 0.393 0.500 0.419 0.388
Autoformer [94] 0.546 0.590 0.565 0.747 0.774 0.918 0.683 0.825
Crossformer [95] 0.500 0.594 0.571 0.775 0.924 0.984 0.737 0.913
TimesNet [96] 0.494 0.594 0.538 0.756 0.794 0.867 0.765 0.944
iTransformer [97] 0.541 0.650 0.600 0.783 0.861 0.931 0.791 0.963
TSMixer [98] 0.488 0.534 0.465 0.689 0.770 0.797 0.808 0.931
TimeMixer [99] 0.577 0.658 0.571 0.776 0.822 0.887 0.824 0.935
FreTS [100] 0.623 0.688 0.546 0.737 0.887 0.950 0.751 0.762
PatchTST [5] 0.592 0.675 0.604 0.795 0.841 0.934 0.695 0.928
LLMTime [101] 0.587 0.657 0.519 0.643 0.802 0.817 0.769 0.803
PromptCast [79] 0.499 0.365 0.418 0.607 0.727 0.768 0.696 0.871
OFA [66] 0.501 0.606 0.512 0.745 0.774 0.879 0.851 0.977
FSCA [73] 0.563 0.647 0.592 0.790 0.820 0.891 0.872 0.977
Time-LLM [69] 0.613 0.699 0.589 0.792 0.671 0.864 0.733 0.912
TimeCMA [72] 0.636 0.731 0.559 0.727 0.729 0.828 0.693 0.843
MM-iTransformer [47] 0.608 0.689 0.605 0.793 0.926 0.986 0.901 0.990
MM-PatchTST [47] 0.621 0.718 0.619 0.812 0.863 0.968 0.780 0.929
TimeCAP [81] 0.668 0.742 0.611 0.801 0.954 0.983 0.942 0.988

### v3

# TimeXL: Explainable Multi-modal Time Series Prediction with LLM-in-the-Loop

@@ 1 Introduction
In the modern big-data era, time series analysis has become indispensable for understanding real-world system behaviors and guiding downstream decision-making tasks across numerous domains, including healthcare, traffic, finance, and weather jin2018treatment ; guo2019attention ; zhang2017stock ; qin2017dual . Although deep learning models have demonstrated success in capturing complex temporal dependencies nietime ; deng2021graph ; zhang2022self ; liu2024koopa , real-world time series are frequently influenced by external information beyond purely temporal factors. Such additional context, which may come from textual narratives (e.g., finance news dong2024fnspid or medical reports king2023multimodal ), can offer critical insights for more accurate forecasting and explainability.
Recent multi-modal approaches for time series have shown promise by integrating rich contextual signals from disparate data sources, thereby improving performance across a range of tasks including forecasting, classification, imputation, and retrieval niu2023deep ; leemoat ; zhao2022vimts ; bamford2023multi . While these approaches utilize supplementary data to enhance predictive accuracy, they often lack explicit mechanisms to systematically reason and explain about why or how contextual signals affect outcomes. This gap in interpretability poses significant barriers for high-stakes applications such as finance and healthcare, where trust and transparency are paramount.
Meanwhile, Large Language Models (LLMs) achiam2023gpt ; team2023gemini have risen to prominence for their remarkable ability to process and reason over textual data across domains, enabling tasks like sentiment analysis, question answering, and content generation in zero- and few-shot settings zhang2024sentiment ; kamalloo2023evaluating ; wang2024c3llm . Recent advancements in agentic designs further augment LLM capabilities, facilitating structured reasoning and iterative decision-making for context-rich scenarios shinn2024reflexion ; madaan2024self . Their encoded domain knowledge makes them natural candidates for supporting real-world multi-modal time series analyses, where textual context (e.g., news or expert notes) plays a vital role liularge ; koa2024learning ; wang2023would .
@@
Motivated by these observations, we introduce TimeXL, a novel framework that adopts a closed-loop workflow of prediction, critique (reflect), and refinement, and unifies a prototype-driven time series encoder with LLM-based reasoning to deliver both accurate and interpretable multi-modal forecasting (Figure 1). Our approach first employs a multi-modal prototype-based encoder to generate preliminary time series predictions alongside human-readable explanations, leveraging case-based reasoning ming2019interpretable ; ni2021interpreting ; jiang2023fedskill from both the temporal and textual modalities. These explanations not only justify the encoder’s predictions but also serve as auxiliary signals to guide an LLM-powered component that further refines the forecasts and contextual rationales.
@@ 2 Related Work
Our work is primarily related to three key areas of time series research: multi-modality, explanation, and the use of LLMs. A more detailed discussion of related methods is provided in Appendix H.
@@
Multi-modal Time Series Analysis. In recent years, multi-modal time series analysis has gained increasing attention in diverse domains such as finance, healthcare and environmental sciences skenderi2024well ; niu2023deep ; zhao2022vimts . Multiple approaches have been proposed to model interactions across different modalities, including fusion and alignment through concatenation li2025languageflowtimetimeseriespaired , attention mechanism leemoat ; chattopadhyay2024context ; su2025multimodal , gating functions zhong2025time , and contrastive learning zheng2024mulan ; li2024frozen . These methods facilitate tasks beyond forecasting and classification, such as retrieval bamford2023multi , imputation zhao2022vimts , and causal discovery zheng2024mulan . Recent research efforts have also advanced the field through comprehensive benchmarks and studies liu2024time ; williams2024context ; jiang2025multi , highlighting the incorporation of new modalities to boost task performance. Meanwhile, others explore transforming time series to other modalities li2023time ; hoo2025tabular ; ni2025harnessing ; xu2025beyond ; chen2024visionts or generating synthetic ones zhong2025time ; liu2025empowering for effective analysis. Nevertheless, these studies primarily focus on improving numerical performance while the deeper and tractable rationale behind why or how the textual or other contextual signals influence time series outcomes remains underexplored.
@@
Time Series Explanation. Recent studies have explored diverse paradigms for time series interpretability. Gradient-based and perturbation-based saliency methods, for example, highlight important features at different time steps ismail2020benchmarking ; tonekaboni2020went , where some works explicitly incorporate temporal structures into models and objectives leungtemporal ; crabbe2021explaining . Surrogate approaches also offer global or local explanations, such as applying Shapley values to time series bento2021timeshap , enforcing model consistency via self-supervised objectives queen2024encoding , or using information-theoretic strategies for coherent explanations liutimex++ . In contrast to saliency or surrogate-based explanations, we adopt a case-based reasoning paradigm ming2019interpretable ; ni2021interpreting ; jiang2023fedskill , which end-to-end generates predictions and built-in explanations from learned prototypes. Our work extends this approach to multi-modal time series by producing human-readable reasoning artifacts for both the temporal and contextual modalities.
@@
LLMs for Time Series Analysis. The rapid development of LLMs achiam2023gpt ; team2023gemini has begun to inspire new directions in time series research ijcai2024p895 ; liang2024foundation . Many existing methods fine-tune pre-trained LLMs on time series tasks and achieves promising results zhou2023one ; bianmulti ; ansari2024chronos , where textual data (e.g., domain instructions, metadata, summaries) are often encoded as prefix embeddings to enrich time series representations jintime ; liu2024timecma ; hu2025contextalignment ; liu2025efficient . These techniques also contribute to the emergence of time series foundation models ansari2024chronos ; das2023decoder ; woo2024unified ; liu2024timer . An alternative line of research leverages the zero-shot or few-shot capabilities of LLMs by directly prompting pre-trained language models with text-converted time series xue2023promptcast or context-laden prompts representing domain knowledge wang2023would , often yielding surprisingly strong performance in real-world scenarios. Furthermore, LLMs can act as knowledge inference modules, synthesizing high-level patterns or explanations that augment standard time series pipelines chen2023chatgpt ; lee2024timecap ; wang2024news . Building on this trend, recent works have explored time series reasoning with LLMs by framing tasks as natural-language problems, including time series understanding & question answering chen2025mtbench ; kong2025position ; kong2025time ; wang2024chattime ; xie2024chatts ; cai2024timeseriesexam , and language-guided inference chen2025mtbench ; kong2025time . Compared with existing works, our work provides a unique synergy between an explainable model and LLMs, enhancing explainable prediction through iterative, grounded, and reflective interaction.
@@ 3.1 Problem Statement
In this paper, we consider a multi-modal time series prediction problem. Each instance is represented by the multi-modal input $(\boldsymbol{x},\boldsymbol{s})$, where $\boldsymbol{x}=(x_{1},x_{2},\cdots,x_{T})\in\mathbb{R}^{N\times T}$ denotes time series data with $N$ variables and $T$ historical time steps, and $\boldsymbol{s}$ denotes the corresponding text data describing the real-world context. The text data $\boldsymbol{s}$ can be further divided into $L$ meaningful segments. Based on the historical time series and textual context, our objective is to predict the future outcome $\boldsymbol{y}$, either as a discrete value for classification tasks, or as a continuous value for regression tasks. We focus on the classification task to reflect discrete decision-making scenarios common in real-world multi-modal applications, where interpretability is often essential for reliable decision support. We also include an evaluation of the regression setting in the Appendix G. There are four major components in the proposed TimeXL framework, a multi-modal prototype encoder $\mathcal{M}_{\mathrm{enc}}$ that provides an initial prediction and case-based explanations, a prediction LLM $\mathcal{M}_{\mathrm{pred}}$ that provides a prediction based on contextual understanding with explanations, a reflection LLM $\mathcal{M}_{\mathrm{refl}}$ that generates feedback, and a refinement LLM $\mathcal{M}_{\mathrm{refine}}$ that refines the textual context based on the feedback. Below, we introduce each component and how they synergize toward better prediction and explanation.
@@ 3.2.1 Multi-modal Sequence Modeling with Prototypes
Sequence Encoder. To capture both temporal and semantic dependencies, we adopt separate encoders for time series ($\mathcal{E}_{\theta}$) and text ($\mathcal{E}_{\phi}$). For $\boldsymbol{x}\in\mathbb{R}^{N\times T}$, the time series encoder $\mathcal{E}_{\theta}$ maps the entire sequence into one or multiple representations, which serve as candidates for prototype learning. Simultaneously, the text input $\boldsymbol{s}$ is first transformed by a frozen pre-trained language model, $\mathrm{PLM}$ (e.g., BERT kenton2019bert or Sentence-BERT reimers-2019-sentence-bert ), to produce embeddings $\boldsymbol{e}_{s}\in\mathbb{R}^{d_{s}\times L}$. These embeddings are then processed by a separate encoder $\mathcal{E}_{\phi}$ to extract meaningful text features. It is worth noting that the choice of $\mathcal{E}_{\theta}$ and $\mathcal{E}_{\phi}$ also affects the granularity of explanations. As we will introduce shortly, the prototypes are learned based on sequence representations and are associated with the counterparts in the input space, where the correspondences are determined by the encoders. In this paper, we choose convolution-based encoders for both modalities to capture the fine-grained sub-sequence (i.e., segment) patterns:
@@ 3.3.1 Model Synergy for Augmented Prediction
Note that $i,j,c$ denotes the prototype index, segment index, and class index, respectively. As $\mathbf{expl}_{\boldsymbol{s}}$ can contain relevant contextual guidance across multiple classes, it augments the input space and removes semantic ambiguity for prediction agent $\mathcal{M}_{\mathrm{pred}}$. Therefore, the prediction is drawn as $\boldsymbol{\hat{y}}_{\mathrm{LLM}}=\mathcal{M}_{\mathrm{pred}}(\boldsymbol{s},\mathbf{expl}_{\boldsymbol{s}})$. The prompt for querying $\mathcal{M}_{\mathrm{pred}}$ is provided in Appendix E, Figure 14.
@@ 3.3.2 Iterative Context Refinement via Reflective Feedback
Given the prediction $\boldsymbol{\hat{y}}_{\mathrm{LLM}}$ generated by the prediction agent $\mathcal{M}_{\mathrm{pred}}$, the reflection agent $\mathcal{M}_{\mathrm{refl}}$ aims to understand the reasoning behind the implicit prediction logic of $\mathcal{M}_{\mathrm{pred}}$. Specifically, it generates a reflective feedback, $\mathrm{Refl}$, by analyzing the input text $\boldsymbol{s}$ and its prediction $\boldsymbol{\hat{y}}_{\mathrm{LLM}}$, against the ground truth $\boldsymbol{y}$, to provide actionable insights for refinement, i.e., $\mathrm{Refl}=\mathcal{M}_{\mathrm{refl}}(\boldsymbol{y},\boldsymbol{\hat{y}}_{\mathrm{LLM}},\boldsymbol{s})$. Guided by the feedback, the refinement agent $\mathcal{M}_{\mathrm{refine}}$ refines the previous text $\boldsymbol{s}_{i}$ into $\boldsymbol{s}_{i+1}$ by selecting and emphasizing the most relevant content, ensuring that important patterns are appropriately contextualized, which is similar to how a domain expert would perform, i.e., $\boldsymbol{s_{i+1}}=\mathcal{M}_{\mathrm{refine}}(\mathrm{Refl},\boldsymbol{s}_{i})$. The prompts for querying $\mathcal{M}_{\mathrm{refl}}$ and $\mathcal{M}_{\mathrm{refine}}$ are provided in Figures 1617 1819, and discussed in Appendix E.
@@
Inputs: Multi-modal time series $(\boldsymbol{x},\boldsymbol{s},\boldsymbol{y})$ with $\mathcal{D}_{\mathrm{train}}$, $\mathcal{D}_{\mathrm{val}}$, $\mathcal{D}_{\mathrm{test}}$, prototype-based encoder $\mathcal{M}_{\mathrm{enc}}$, prediction agent $\mathcal{M}_{\mathrm{pred}}$, reflection agent $\mathcal{M}_{\mathrm{refl}}$, refinement agent $\mathcal{M}_{\mathrm{refine}}$, fusion parameter $\alpha$, max iteration $\tau$, improvement check $\mathrm{Eval(\cdot)}$
Training & Validation:
@@ 4 Experiments
DLinear zeng2023transformers 0.540 0.660 0.255 0.485 0.393 0.500 0.419 0.388
Autoformer wu2021autoformer 0.546 0.590 0.565 0.747 0.774 0.918 0.683 0.825
Crossformer zhang2023crossformer 0.500 0.594 0.571 0.775 0.924 0.984 0.737 0.913
TimesNet wutimesnet 0.494 0.594 0.538 0.756 0.794 0.867 0.765 0.944
iTransformer liuitransformer 0.541 0.650 0.600 0.783 0.861 0.931 0.791 0.963
TSMixer chentsmixer 0.488 0.534 0.465 0.689 0.770 0.797 0.808 0.931
TimeMixer wangtimemixer 0.577 0.658 0.571 0.776 0.822 0.887 0.824 0.935
FreTS yi2024frequency 0.623 0.688 0.546 0.737 0.887 0.950 0.751 0.762
PatchTST nietime 0.592 0.675 0.604 0.795 0.841 0.934 0.695 0.928
LLMTime gruver2023llmtime 0.587 0.657 0.519 0.643 0.802 0.817 0.769 0.803
PromptCast xue2023promptcast 0.499 0.365 0.418 0.607 0.727 0.768 0.696 0.871
OFA zhou2023one 0.501 0.606 0.512 0.745 0.774 0.879 0.851 0.977
FSCA hu2025contextalignment 0.563 0.647 0.592 0.790 0.820 0.891 0.872 0.977
Time-LLM jintime 0.613 0.699 0.589 0.792 0.671 0.864 0.733 0.912
TimeCMA liu2024timecma 0.636 0.731 0.559 0.727 0.729 0.828 0.693 0.843
MM-iTransformer liu2024time 0.608 0.689 0.605 0.793 0.926 0.986 0.901 0.990
MM-PatchTST liu2024time 0.621 0.718 0.619 0.812 0.863 0.968 0.780 0.929
TimeCAP lee2024timecap 0.668 0.742 0.611 0.801 0.954 0.983 0.942 0.988